Nicole has one less than one-fifth as many dimes as quarters. She has $8.00 worth of coins. How many of each coin does she have?
Answer by mananth(16949)

quarters =x
dimes= (1/5)x-1
she has 800 cents
25x+10((1/5)x-1) = 800
25x+2x-10=800
27x=810
/27
x=30 quarters
dimes = (1/5)*30 -1
dimes= 5
CHECK
5*10 + 30*25
50+750=800